# Global bifurcation of localised 2D patterns emerging from spatial heterogeneity

### v1

@@ 1 Introduction
The emergence of patterns with spatially localised extent are important for understanding phenomena in a wide range of applications from the formation of tornadoes [53, 15], patches and fairy circles in dryland vegetation [12, 39, 33, 52], and spikes on the surface of a magnetic fluid [50, 59]; see [7] for a review of localised pattern formation. The key bifurcation of these localised patterns from quiescence is due to a pattern-forming instability [7]. While there is extensive theory for the emergence of one-dimensional localised patterns, to date there remains no existence proof for fully localised 2D patterns emerging from a Turing instability beyond the axisymmetric case. Fundamentally, this is due to the fact that the kernel of the linear operator at a 1D pattern-forming instability is finite dimensional but in 2D it is infinite-dimensional. We note that there have been recent advances in approximating these patterns [34, 35, 36], as well as in proving their existence at certain parameter regimes [61, 14, 51] via rigorous numerics.
@@
The Crandall–Rabinowitz theorem allows one to prove the existence of small amplitude solutions bifurcating from a given base state while analytic global bifurcation theory—originally due to Dancer [20] and later developed by Buffoni and Toland [11] and Constantin, Strauss and Vărvărucă [18]—allows one to continue small amplitude solutions to large amplitude. The Crandall–Rabinowitz theorem is typically applied to bounded domains or spatially periodic solutions; while localised solutions have been considered in some cases [64, 60], the application of Crandall–Rabinowitz theory is more difficult due to a lack of standard embedding and compactness properties in infinitely-extended spatial domains. While analytic global bifurcation theory has become invaluable in proving the existence of large amplitude water waves, including some recent applications to solitary waves (see [31, Section 5.6] and references therein), very few results exist in the pattern formation literature (see for instance [54, 24, 4, 25] which use the global bifurcation theory of Rabinowitz [57], and the more recent work of [10, 13] which use analytic global bifurcation theory). A key difficulty with applying the theory of Buffoni and Toland [11] to localised solutions is a possible lack of compactness of solution sets. While in bounded domains (or, equivalently, for periodic solutions) local compactness of the solution set is guaranteed by a combination of Schauder estimates and Rellich–Kondrachev-type embeddings, these results no longer hold in infinitely extended domains. A crucial change in this approach was provided by Chen, Walsh and Wheeler in [17], where the authors adapted the analytic global continuation theorem of Buffoni and Toland so that local compactness was no longer an assumption, but rather a property of the solution set that could fail. To the best of the authors’ knowledge, there are no global bifurcation results for localised patterns in pattern-forming systems, and so this paper presents a new application of the theory for 2D fully localised patterns with dihedral symmetry.
@@
$\partial_{t}u=\mathcal{L}_{\varepsilon}(\Delta)u+V_{\varepsilon}(|\mathbf{x}|)\,u+\mathcal{N}_{\varepsilon}(u),\qquad\qquad u\in H^{2m}(\mathbb{R}^{d};\mathbb{R}^{N}),\qquad\varepsilon\in\Lambda\subseteq\mathbb{R},$ (1.1)
with $\mathcal{L}_{\varepsilon}(\Delta):H^{2m}(\mathbb{R}^{d};\mathbb{R}^{N})\to L^{2}(\mathbb{R}^{d};\mathbb{R}^{N})$ an isotropic linear operator (where $\Delta$ denotes the $d$-dimensional Laplace operator and $m>d/4$), $V_{\varepsilon}$ a localised radially-symmetric potential, and $\mathcal{N}_{\varepsilon}$ a real-analytic function with $\mathcal{N}_{\varepsilon}(0)=\mathcal{N}_{\varepsilon}^{\prime}(0)=0$ for all $\varepsilon\in\Lambda$. We additionally assume that $\mathcal{L}_{\varepsilon},V_{\varepsilon},\mathcal{N}_{\varepsilon}$ are each analytic in $\varepsilon\in\Lambda$, and that the following hypotheses are satisfied.
@@
    there exists some $\varepsilon_{0}\in\Lambda$ such that $\mathrm{ker}\,(\mathcal{L}_{\varepsilon_{0}}(\Delta)+V_{\varepsilon_{0}})=\mathrm{span}\,\{v_{0}\}$ for some $v_{0}\in X\subseteq H^{2m}(\mathbb{R}^{d};\mathbb{R}^{N})$ [^1] One must restrict to some subspace $X$ of $H^{2m}(\mathbb{R}^{d};\mathbb{R}^{N})$ in order to remove innate symmetries of the problem and obtain a one-dimensional kernel.; and
@@
Hypothesis (i) supposes that the operator $\mathcal{L}_{\varepsilon}(\Delta)$ and the linearisation of (1.1) about some $u=\phi\in H^{2m}(\mathbb{R}^{d};\mathbb{R}^{N})$ are Fredholm operators with index zero. In particular, the essential spectrum of each operator is bounded away from zero and so any changes in the linear stability of solutions to (1.1) is driven entirely by eigenvalues crossing the imaginary axis. Hypothesis (ii) supposes that, for some parameter value $\varepsilon=\varepsilon_{0}$, there is one eigenvalue lying on the imaginary axis with a single corresponding eigenfunction $v_{0}$. Finally, hypothesis (iii) supposes that the equation (1.1) has nondegenerate dependence on the parameter $\varepsilon$ at $\varepsilon=\varepsilon_{0}$.
@@
Here, $\mathcal{E}(0)=\varepsilon_{0}$, $v(0)=0$, $v^{\prime}(0)=v_{0}$, and the curve $\mathcal{C}$ possesses a local analytic parametrisation around each of its points. The local and global bifurcation results are proven in an open subspace $(\varepsilon,u)\in U\subseteq\Lambda\times H^{2m}(\mathbb{R}^{d};\mathbb{R}^{N})$, and so a solution $u=v(s)$ of (1.1) will remain spatially localised along the curve $\mathcal{C}$.
@@
where $u=u(t,\mathbf{x})\in\mathbb{R}$, $\mathbf{x}\in\mathbb{R}^{2}$, $R>0$ denotes the width of the potential $V$, and $f$ is real-analytic. This equation naturally destabilises the trivial state in a finite circle of radius $R$ while outside this region the trivial state is linearly stable. We seek steady solutions $u=u(\mathbf{x})$ in $H^{4}_{\mathrm{e}}$—the subspace of functions in $H^{4}(\mathbb{R}^{2})$ which are also even in $y$, i.e. u(x,y)=u(x,y)u(x,y)=u(x,-y)—where we solve the steady equation
@@
Our aim is then to prove the schematic picture shown in Figure 2(a). We note that we prove the existence of fully localised patterns with $\mathbb{D}_{k}$ symmetry for any $k\in\mathbb{N}_{0}$, and not just the $\mathbb{D}_{0}$, $\mathbb{D}_{1}$ and $\mathbb{D}_{6}$ localised patterns shown in Figure 2(a). We employ analytic bifurcation theory in order to prove the existence of the bifurcating localised patterns, as illustrated in Figure 2(b); we first prove the existence of a unique non-trivial solution branch in a neighbourhood of the bifurcation point (Figure 2($b_{1}$)), which we call the local branch, before extending solutions to large amplitude (Figure 2($b_{2}$)), which we call the global branch. The global solution branch is either unbounded, approaches the $\varepsilon=0$ axis, forms a closed loop, or loses compactness, as shown in Figure 2(ii)-(iviv).
@@ 3.1 Local bifurcation
We begin by proving the existence of a unique bifurcation branch in a neighbourhood of the point $(\varepsilon,u)=(\varepsilon_{k,n},0)$, which we call the local bifurcation branch. We will use an analytic version of the Crandall–Rabinowitz theorem (cf. Buffoni and Toland [11, Theorem 8.3.1]), which we state as Theorem A.1 in the appendix. In order to prove that a local bifurcation occurs, we need to verify the following conditions:
@@ 3. $(iii)$
and $\mathrm{ker}(\mathcal{L})=\mathrm{span}(v_{0})$. There exists a projection $\mathcal{P}$ with $\mathrm{Im}(\mathcal{P})=\mathrm{ker}(\mathcal{L})$ given by
@@
and so $\mathcal{P}\left(\mathrm{d}_{1}\mathrm{d}_{2}\mathscr{F}[\varepsilon_{0},0](1,v_{0})\right)=0$ if and only if $v_{0}\,(1+\Delta)v_{0}=0$. For any function $v\in H^{1}(\mathbb{R}^{n};\mathbb{R})$ satisfying $(1+\Delta)v=0$, we note that Pokhozhaev’s identity (cf. Berestycki and Lions [6, (2.1)])
@@
Utilising the analytic properties of the local bifurcation curve, we also apply a supplementary theorem presented in Groves and Horn [30, Theorem 4.1] to characterise the local bifurcation curve $\mathcal{C}_{\delta}$. We state this result as Theorem A.2 in the appendix, which we apply in the following corollary.
@@
Fix $R>0$, $k\in\mathbb{N}_{0}$ and $n\in\mathbb{N}$ as in Theorem 3.3, so that a local bifurcation occurs at $\varepsilon=\varepsilon_{k,n}$, with bifurcating profile $u(\mathbf{x})=s\,u_{k,n}(\mathbf{x})+\mathcal{O}(s^{2})$ for $|s|<\delta$. The local bifurcation is a pitchfork for all $k\neq 0$; for $k=0$ the bifurcation is a pitchfork if $f^{\prime\prime}(0)\,\langle u^{\vphantom{2}}_{k,n},u_{k,n}^{2}\rangle_{L^{2}}=0$ and a transcritical bifurcation otherwise.
@@
The characterisation of the local bifurcation curve follows directly from Theorem A.2, where we note that
$\varepsilon_{1}=-\frac{\mathcal{P}\left(\frac{1}{2!}\,\mathrm{d}_{2}^{2}\mathscr{F}[\varepsilon_{k,n},0](u_{k,n},u_{k,n})\right)}{\mathcal{P}\left(\mathrm{d}_{1}\mathrm{d}_{2}\mathscr{F}[\varepsilon_{k,n},0](1,u_{k,n})\right)}=-\frac{\varepsilon_{k,n}}{2}\,f^{\prime\prime}(0)\,\frac{\langle u^{\vphantom{2}}_{k,n},u_{k,n}^{2}\rangle_{L^{2}}}{\|(1+\Delta)u_{k,n}\|^{2}_{L^{2}}}.$
In particular, for $k\neq 0$ we obtain $u_{k,n}^{2}(r\cos(\theta),r\sin(\theta))=\frac{1}{2}v_{k,n}^{2}(r)+\frac{1}{2}v_{k,n}^{2}(r)\,\cos(2k\theta)$ and so, by the orthogonality of the angular Fourier modes, it follows that
$\langle u^{\vphantom{2}}_{k,n},u_{k,n}^{2}\rangle_{L^{2}}=0$
@@
We only consider the primary bifurcation branch where the trivial state first destabilises; the eigenvalues of the bifurcating solution will be a small perturbation of the eigenvalues of the trivial state (cf. Kato [47, Theorem VIII-2.6]), and so each bifurcating solution will be unstable for an unstable trivial state.
@@
Since we have assumed that the trivial state destabilises at the bifurcation point, it follows that $\lambda^{*}|_{s=0}=0$. We note that $\lambda^{*}|_{s=0}=0$ is simple, by Hypothesis 1, and so there exists an analytic function $\Lambda:(-\delta,\delta)\to\mathbb{R}$ such that $\lambda^{*}=s\,\Lambda(s)$ in a neighbourhood of the curve $\mathcal{C}_{\delta}$ (see Kato [47, VII-§3.2]). We consider the following Taylor expansions
@@
For particular choices of the nonlinearity $f(u)$, we can immediately predict the bifurcation structure of these localised patterns. In particular, for $f(u)=-u^{3}$ (or, indeed, any odd analytic function $f$ with $f^{\prime\prime\prime}(0)<0$) every bifurcation will be a supercritical pitchfork, and so the primary bifurcating branches will both be stable.

### v2

@@ 1 Introduction
The emergence of patterns with spatially localised extent are important for understanding phenomena in a wide range of applications from the formation of tornadoes [62, 17], patches and fairy circles in dryland vegetation [15, 43, 41, 59], and spikes on the surface of a magnetic fluid [57, 68]; see [10] for a review of localised pattern formation. The key bifurcation of these localised patterns from quiescence is due to a pattern-forming instability [10]. While there is extensive theory for the emergence of one-dimensional localised patterns, to date there remains no existence proof for fully localised 2D patterns emerging from a Turing instability beyond the axisymmetric case [56, 60, 61, 42]. Fundamentally, this is due to the fact that the kernel of the linear operator at a 1D pattern-forming instability is finite dimensional but in 2D it is infinite-dimensional. We note that there have been recent advances in approximating these patterns [38, 39, 40], as well as in proving their existence at certain parameter regimes [74, 16, 58] via rigorous numerics.
@@
The Crandall–Rabinowitz theorem allows one to prove the existence of small amplitude solutions bifurcating from a given base state while analytic global bifurcation theory—originally due to Dancer [22] and later developed by Buffoni and Toland [14] and Constantin, Strauss and Vărvărucă [20]—allows one to continue small amplitude solutions to large amplitude. The Crandall–Rabinowitz theorem is typically applied to bounded domains or spatially periodic solutions; while localised solutions have been considered in some cases [78, 73], the application of Crandall–Rabinowitz theory is more difficult due to a lack of standard embedding and compactness properties in infinitely-extended spatial domains. While analytic global bifurcation theory has become invaluable in proving the existence of large amplitude water waves, including some recent applications to solitary waves (see [34, Section 5.6] and references therein), very few results exist in the pattern formation literature (see for instance [63, 27, 5, 28] which use the global bifurcation theory of Rabinowitz [66], and the more recent work of [13, 9] which use analytic global bifurcation theory). A key difficulty with applying the theory of Buffoni and Toland [14] to localised solutions is a possible lack of compactness of solution sets. While in bounded domains (or, equivalently, for periodic solutions) local compactness of the solution set is guaranteed by a combination of Schauder estimates and Rellich–Kondrachev-type embeddings, these results no longer hold in infinitely extended domains. A crucial change in this approach was provided by Chen, Walsh and Wheeler in [19], where the authors adapted the analytic global continuation theorem of Buffoni and Toland so that local compactness was no longer an assumption, but rather a property of the solution set that could fail. To the best of the authors’ knowledge, there are no global bifurcation results for localised patterns in pattern-forming systems, and so this paper presents a new application of the theory for 2D fully localised patterns with dihedral symmetry.
@@
$\partial_{t}u=\mathcal{L}_{\varepsilon}(\Delta)u+V_{\varepsilon}(\mathbf{x})\,u+\mathcal{N}_{\varepsilon}(u),\qquad\qquad u\in H^{2m}(\mathbb{R}^{d};\mathbb{R}^{N}),\qquad\varepsilon\in\Lambda\subseteq\mathbb{R},$ (1.1)
with $\mathcal{L}_{\varepsilon}(\Delta):H^{2m}(\mathbb{R}^{d};\mathbb{R}^{N})\to L^{2}(\mathbb{R}^{d};\mathbb{R}^{N})$ an isotropic linear operator (where $\mathcal{L}_{\varepsilon}(\cdot)$ is given by an $m$’th order polynomial, $\Delta$ denotes the $d$-dimensional Laplace operator, and $m>d/4$), $V_{\varepsilon}$ a localised potential, and $\mathcal{N}_{\varepsilon}$ a real-analytic function with $\mathcal{N}_{\varepsilon}(0)=\mathcal{N}_{\varepsilon}^{\prime}(0)=0$ for all $\varepsilon\in\Lambda$. We additionally assume that $\mathcal{L}_{\varepsilon},V_{\varepsilon},\mathcal{N}_{\varepsilon}$ are each analytic in $\varepsilon\in\Lambda$, and that the following hypotheses are satisfied.
@@
Let $X\subseteq H^{2m}(\mathbb{R}^{d};\mathbb{R}^{N})$ [^1] One must restrict to some subspace $X$ of $H^{2m}(\mathbb{R}^{d};\mathbb{R}^{N})$ in order to remove innate symmetries of the problem and obtain a one-dimensional kernel in hypothesis $(ii)$. be an invariant subspace of (1.1) and (with slight abuse of notation) replace $\mathcal{L}_{\varepsilon}(\Delta)$ with its restriction onto $X$. We then suppose that
@@
    there exists some $\varepsilon_{0}\in\Lambda$ such that $\mathrm{ker}\,(\mathcal{L}_{\varepsilon_{0}}(\Delta)+V_{\varepsilon_{0}})=\mathrm{span}\,\{v_{0}\}$ with $v_{0}\in X$; and
@@
Hypothesis (i) supposes that the operator $\mathcal{L}_{\varepsilon}(\Delta)$ and the linearisation of (1.1) about some $u=\phi\in H^{2m}(\mathbb{R}^{d};\mathbb{R}^{N})$ are Fredholm operators with index zero. In particular, the essential spectrum of each operator is bounded away from zero and so any change in the linear stability of solutions to (1.1) is driven entirely by eigenvalues crossing the imaginary axis. Hypothesis (ii) supposes that, for some parameter value $\varepsilon=\varepsilon_{0}$, there is one eigenvalue lying on the imaginary axis with a single corresponding eigenfunction $v_{0}$. Finally, hypothesis (iii) supposes that the equation (1.1) has nondegenerate dependence on the parameter $\varepsilon$ at $\varepsilon=\varepsilon_{0}$.
@@
Here, $\mathcal{E}(0)=\varepsilon_{0}$, $v(0)=0$, $v^{\prime}(0)=v_{0}$, and the curve $\mathcal{C}$ possesses a local analytic parametrisation around each of its points. The local and global bifurcation results are proven in an open subspace $(\varepsilon,u)\in U\subseteq\Lambda\times X$, and so a solution $u=v(s)$ of (1.1) will remain spatially localised along the curve $\mathcal{C}$.
@@
where $u=u(t,\mathbf{x})\in\mathbb{R}$, $\mathbf{x}\in\mathbb{R}^{2}$, $R>0$ denotes the width of the potential $V$, and $f$ is real-analytic with $f(0)=f^{\prime}(0)=0$. This equation naturally destabilises the trivial state in a finite circle of radius $R$ while outside this region the trivial state is linearly stable. While (1.2) is a considerably simpler model than the general PDE system (1.1), the theory presented in later sections either follows identically or can be extended to (1.1) (under the assumption that hypotheses (i)-(iii) are satisfied); see Remark 1.3 at the end of this section for further discussion of the generality of this approach.
We seek steady solutions $u=u(\mathbf{x})$ in $H^{4}_{\mathrm{e}}$—the subspace of functions in $H^{4}(\mathbb{R}^{2})$ which are also even in $y$, i.e. u(x,y)=u(x,y)u(x,y)=u(x,-y)—where we solve the steady equation
@@
Our aim is then to prove the schematic picture shown in Figure 2(a). We note that we prove the existence of fully localised patterns with $\mathbb{D}_{k}$ symmetry for any $k\in\mathbb{N}_{0}$, and not just the $\mathbb{D}_{0}$, $\mathbb{D}_{1}$ and $\mathbb{D}_{6}$ localised patterns shown in Figure 2(a). We employ analytic bifurcation theory in order to prove the existence of the bifurcating localised patterns, as illustrated in Figure 2(b); we first prove the existence of a unique non-trivial solution branch in a neighbourhood of the bifurcation point (Figure 2($b_{1}$)), which we call the local branch, before extending solutions to large amplitude (Figure 2($b_{2}$)), which we call the global branch. The global solution branch is either unbounded, approaches the $\varepsilon=0$ axis, forms a closed loop, or loses compactness, as shown in Figure 2(ii)-(iviv). In the case of radially symmetric patterns, we are able to prove that the global solution branch never loses compactness; an equivalent result for non-radially symmetric patterns remains an open problem.
@@
We briefly discuss how Results 1 and 2—proven for the Swift–Hohenberg equation (1.2)—can be extended to more general PDE systems of the form (1.1).
Remark 1.3.
Suppose we consider fully localised solutions to (1.1), where we assume
  • $\mathcal{L}_{\varepsilon}(\cdot)$ is given by an $m$’th order polynomial, so that $\mathcal{L}_{\varepsilon}(\Delta):H^{2m}(\mathbb{R}^{d};\mathbb{R}^{N})\to L^{2}(\mathbb{R}^{d};\mathbb{R}^{N})$;
  • $\mathcal{N}_{\varepsilon}$ is a real-analytic function with $\mathcal{N}_{\varepsilon}(0)=\mathcal{N}_{\varepsilon}^{\prime}(0)=0$ for all $\varepsilon\in\Lambda$, so that the linearisation of (1.1) is given by $(\mathcal{L}_{\varepsilon}(\Delta)+V_{\varepsilon})\,u$;
  • $m>d/4$, so that $H^{2m}(\mathbb{R}^{d};\mathbb{R}^{N})$ is a Banach algebra and hence $\mathcal{N}_{\varepsilon}:H^{2m}(\mathbb{R}^{d};\mathbb{R}^{N})\to L^{2}(\mathbb{R}^{d};\mathbb{R}^{N})$ analytically;
  • $V_{\varepsilon}$ is chosen so that $\lim_{|\mathbf{x}|\to\infty}V_{\varepsilon}(\mathbf{x})=0$, and the multiplication operator $u\mapsto V_{\varepsilon}\,u$ maps from $H^{2m}(\mathbb{R}^{d};\mathbb{R}^{N})$ to $L^{2}(\mathbb{R}^{d};\mathbb{R}^{N})$; and
  • $\mathcal{L}_{\varepsilon},V_{\varepsilon},\mathcal{N}_{\varepsilon}$ are each analytic in $\varepsilon\in\Lambda$.
Then, hypotheses (i-iii) guarantee that (1.1) satisfies the conditions for a Crandall-Rabinowitz local bifurcation. Since any solution $\phi\in H^{2m}(\mathbb{R}^{d};\mathbb{R}^{N})$ decays to zero as $|\mathbf{x}|\to\infty$, the linearisation about $\phi$ possesses the same Fredholm properties as $\mathcal{L}_{\varepsilon}(\Delta)$ which, coupled with the analytic dependence of (1.1) on $(\varepsilon,u)$, allows one to also globally continue the local bifurcation curves to large amplitude. For radially-symmetric patterns, our proof that global branches do not lose compactness also applies directly to (1.1) under the assumptions above.
The key difficulty in considering the general PDE system (1.1) is verifying hypotheses (ii) and (iii). Our choice of a piecewise-constant, radially-symmetric potential $V$ in (1.2) allows for the explicit analysis of the point spectrum of $\mathcal{L}$ in Lemma 2.4; even extending our analysis to a potential without radial symmetry results in a significantly more difficult eigenvalue problem, which may not be analytically tractable.
@@ 3.1 Local bifurcation
We begin by proving the existence of a unique bifurcation branch in a neighbourhood of the point $(\varepsilon,u)=(\varepsilon_{k,n},0)$, which we call the local bifurcation branch. We will use an analytic version of the Crandall–Rabinowitz theorem (cf. Buffoni and Toland [14, Theorem 8.3.1]), which we state as Theorem A.1 in the appendix. In order to prove that a local bifurcation occurs, we need to verify the following conditions:
@@ 3. $(iii)$
and $\mathrm{ker}(\mathcal{L})=\mathrm{span}(v_{0})$. There exists a projection $\mathcal{P}$ with $\mathrm{Im}(\mathcal{L})=\mathrm{ker}(\mathcal{P})$ given by
@@
and so $\mathcal{P}\left(\mathrm{d}_{1}\mathrm{d}_{2}\mathscr{F}[\varepsilon_{0},0](1,v_{0})\right)=0$ if and only if $v_{0}\,(1+\Delta)v_{0}=0$. For any function $v\in H^{1}(\mathbb{R}^{n};\mathbb{R})$ satisfying $(1+\Delta)v=0$, we note that Pokhozhaev’s identity (cf. Berestycki and Lions [8, (2.1)])
@@
Utilising the analytic properties of the local bifurcation curve, we also apply a supplementary theorem presented in Groves and Horn [32, Theorem 4.1] to characterise the local bifurcation curve $\mathcal{C}_{\delta}$. We state this result as Theorem A.2 in the appendix, which we apply in the following corollary.
@@
Fix $R>0$, $k\in\mathbb{N}_{0}$ and $n\in\mathbb{N}$ as in Theorem 3.3, so that a local bifurcation occurs at $\varepsilon=\varepsilon_{k,n}$, with Taylor expansions
$\varepsilon=\varepsilon_{k,n}+s\,\varepsilon_{1}+s^{2}\,\varepsilon_{2}+\mathcal{O}(s^{3}),\qquad\qquad u(\mathbf{x})=s\,u_{k,n}(\mathbf{x})+s^{2}\,v_{1}(\mathbf{x})+\mathcal{O}(s^{3})$
for $|s|<\delta$, $\varepsilon_{1},\varepsilon_{2}\in\mathbb{R}$, and $v_{1}\in\ker(\mathcal{Q})$, where $\mathcal{Q}:L^{2}_{\mathrm{e}}\to L^{2}_{\mathrm{e}}$ is a projection with $\mathrm{Im}(\mathcal{Q})=\ker(\mathcal{L})$.
The coefficient $\varepsilon_{1}$ is given by
$\varepsilon_{1}={}$ $-\frac{\varepsilon_{k,n}}{2}\,f^{\prime\prime}(0)\,\frac{\langle u_{k,n},u_{k,n}^{2}\rangle_{L^{2}}}{\|(1+\Delta)\,u_{k,n}\|^{2}_{L^{2}}}.$
and the bifurcation is transcritical if $\varepsilon_{1}\neq 0$. If $\varepsilon_{1}=0$, the coefficient $\varepsilon_{2}$ is given by
$\varepsilon_{2}={}$ $-\frac{\varepsilon_{k,n}}{2}\,f^{\prime\prime}(0)\,\frac{\langle u_{k,n},u_{k,n}\,v_{1}\rangle_{L^{2}}}{\|(1+\Delta)\,u_{k,n}\|_{L^{2}}^{2}}-\frac{\varepsilon_{k,n}}{4}\,\,f^{\prime\prime\prime}(0)\,\frac{\|u_{k,n}\|_{L^{4}}^{4}}{\|(1+\Delta)\,u_{k,n}\|_{L^{2}}^{2}}$
and the bifurcation is a subcritical (supercritical) pitchfork if $\varepsilon_{2}<0$ ($\varepsilon_{2}>0$). In particular, $\varepsilon_{1}=0$ for all $k\neq 0$.
@@
The characterisation of the local bifurcation curve follows directly from Theorem A.2, where we note that for $k\neq 0$ we obtain $u_{k,n}^{2}(r\cos(\theta),r\sin(\theta))=\frac{1}{2}v_{k,n}^{2}(r)+\frac{1}{2}v_{k,n}^{2}(r)\,\cos(2k\theta)$ and so, by the orthogonality of the angular Fourier modes, it follows that
$\varepsilon_{1}={}$ $-\frac{\varepsilon_{k,n}}{2}\,f^{\prime\prime}(0)\,\frac{\langle u_{k,n},u_{k,n}^{2}\rangle_{L^{2}}}{\|(1+\Delta)\,u_{k,n}\|^{2}_{L^{2}}}=0$
@@
We only consider the primary bifurcation branch where the trivial state first destabilises; the eigenvalues of the bifurcating solution will be a small perturbation of the eigenvalues of the trivial state (cf. Kato [51, Theorem VIII-2.6]), and so each bifurcating solution will be unstable for an unstable trivial state.
@@
Since we have assumed that the trivial state destabilises at the bifurcation point, it follows that $\lambda^{*}|_{s=0}=0$. We note that $\lambda^{*}|_{s=0}=0$ is simple, by Hypothesis 1, and so there exists an analytic function $\Lambda:(-\delta,\delta)\to\mathbb{R}$ such that $\lambda^{*}=s\,\Lambda(s)$ in a neighbourhood of the curve $\mathcal{C}_{\delta}$ (see Kato [51, VII-§3.2]). We consider the following Taylor expansions
@@
For particular choices of the nonlinearity $f(u)$, we can immediately predict the bifurcation structure of these localised patterns. In particular, for $f(u)=-u^{3}$ (or, indeed, any analytic function $f$ with $f^{\prime\prime}(0)=0$ and $f^{\prime\prime\prime}(0)<0$) every bifurcation will be a supercritical pitchfork, and so the primary bifurcating branches will both be stable.